Sulfuric Acid is a colorless strong acid that is miscible with water. It is also known as "dipping acid" because it is very corrosive and dissolves most metals. What is the formula mass of this acid if its chemical formula is H₂SO4?

1
Expert's answer
2021-11-11T11:56:01-0500

H "= 1 \times 2 = 2"

S = 32

O "= 16 \times 4 = 64"

The formula mass = 2 + 32 + 64 = 98 g/mol